As per one of the ten principles of economics, a country’s standard of living depends on...

  1. As per one of the ten principles of economics, a country’s standard of living depends on its

    a.

    ability to produce goods and services.

    b.

    quantity of physical capital.

    c.

    ability to thrive economically without having to interact with other countries.

    d.

    abundance of natural resources.

  2. Refer to the information below for the economy of Greenland:

    Real GDP per person in 2018 = $7,250

    Real GDP per Person in 2017 = $7,050

    By what percentage the real GDP per person grow in this economy?

    1.0 percent

    1.5 percent

    2.8 percent

    6.0 percent

  3. Uzbekistan government decided to help the domestic producers to boost their business. For this they removed several restrictions on trade

    This would be an outward oriented policy and economists believe that it will help for the economic growth of Uzbekistan.

    This would be an inward oriented policy and economists believe that it will help for the economic growth of Uzbekistan.

    This would be an outward oriented policy and economists believe that it will cause adverse affect on the economic growth of Uzbekistan.

    This would be an inward oriented policy and economists believe that it will cause adverse affect on the economic growth of Uzbekistan

  4. All else equal, which of the following would tend to cause real GDP per person to rise?

    encouraging inward-oriented trade policies

    an increase in investment in human capital

    a weakening of property rights

    discouraging outward oriented trade policies
Answer #1

1. Option A. ability to produce goods and services.

Explanation: The standard of living of a country depends on the amount of goods or services the country can produce.

2. Option 3. 2.8%

Explanation: The % change in real GDP = (7,250 - 7,050)/7,050 * 100 = 200/7,050 * 100 = 20,000/7,050 = 2.8%

3. Option 1. his would be an outward oriented policy and economists believe that it will help for the economic growth of Uzbekistan.

Explanation: Opening for international trade would be an outward policy. Also, trade helps in economic development.

4. Option 2. An increase in investment in human capital

Explanation: An increase in the human capital would result in higher productivity and an improvement in the standard of living.